As a stockholder in Randolph Corporation, you receive its annual report.
In the financial statements, Randolph has reported that the after-tax (net)
income is $300 million. With 150 million shares of common stock
outstanding, Randolph announced to distribute $100 million of dividends
to its shareholders. The stock is now sold for $20 per share.

i. Calculate the value of investor’s wealth who holds 4,000 shares of Randolph
Corporation before the ex-dividend date,
ii. Calculate the dividend income of an investor who holds 4,000 shares of
Randolph Corporation until the ex-dividend date,
iii. Calculate the value of shareholding on the ex-dividend date of an investor who
holds 4,000 shares of Randolph Corporation.
1
Expert's answer
2019-12-06T09:38:56-0500

Finance

We need to Calculate the value of investor’s wealth, dividend income of an investor and value of shareholding on the ex-dividend date.


Solution:


(i).

Given,

Number of shares = 4000


Sale price of a share =$20


Investor’s wealth who holds 4,000 shares of Randolph Corporation before the ex-dividend date


"= Number \space of \space shares \times sale \space price"

"=4000 \times 20 = \$80000"

(ii).


Dividend income of an investor who holds 4,000 shares of Randolph Corporation until the ex-dividend date


"= Dividend \space per \space share \times number \space of \space shares \space hold"

"= \frac {100} {150} \times 4000 = \$2666.67"


(iii).


The value of shareholding on the ex-dividend date of an investor who holds 4,000 shares of Randolph Corporation.

"= Number \space of \space shares \times \frac {market value }{ Number \space of shares \space of common \space stock}"

"= 4000 \times \frac {3000 million - 100 million} {150 million}"

"= 4000 \times 19.33 = \$ 77333.33"


( Market value of Randolph Corporation before Ex-Dividend Date = $ 20 per share * 150 million shares = $ 3000 million)


Answer: Yes.


Dividend income of an investor who holds 4,000 shares of  Randolph Corporation until the ex-dividend date =$80000


Dividend income of an investor who holds 4,000 shares of Randolph Corporation until the ex-dividend date =$2666.67


The value of shareholding on the ex-dividend date of an investor who holds 4,000 shares of Randolph Corporation= $77333.33